Find the general solution of

(y-xu)p+(x+yu)q=x2+y2

1
Expert's answer
2021-12-07T10:15:32-0500

Given, "(y-xu)p+(x+yu)q = x^{2}+y^{2}".

This equation is of the form "Pp+Qq=R" (Lagrange's linear partial differential equation).

Here, "P = y-xu, Q=(x+yu), R = x^{2}+y^{2}".


The subsidiary equations are

"\dfrac{dx}{P}=\dfrac{dy}{Q}=\dfrac{du}{R}\\
\dfrac{dx}{y+xu}=\dfrac{dy}{-(x+yu)}=\dfrac{du}{x^{2}+y^{2}}~~~~~~~~~~~-(1)".

Each ratio of (1) is equal to "\dfrac{xdx+ydy}{(x^{2}-y^{2})u} = \dfrac{dx+dy}{(1-u)(y-x)}".

Let us consider,

"\dfrac{xdx+ydy}{(x^{2}-y^{2})u} = \dfrac{du}{x^{2}-y^{2}}\\
xdx+ydy=udu\\
\text{Integrating, }\\ \int xdx + \int ydy = \int zdz\\
\dfrac{x^{2}}{2}+\dfrac{y^{2}}{2} =\dfrac{z^{2}}{2}+\dfrac{c_{1}}{2}\\
x^{2}+y^{2}-u^{2}=c_{1}"


Consider,

"\dfrac{dx+dy}{(1-u)(y-x)}=\dfrac{du}{(x+y)(x-y)}\\
\dfrac{dx+dy}{1-u}=\dfrac{du}{-(x+y)}\\
-(x+y)d(x+y)=(1-u)du\\
\text{Integrating,}\\
-\int (x+y)d(x+y)=\int (1+z)dz\\
-\dfrac{(x+y)^2}{2}=\dfrac{(1-z)^{2}}{-2}+\dfrac{c_{2}}{2}\\
(1-u)^{2}-(x+y)^{2}=c_{2}"


The general solution is,

"\phi(c_{1},c_{2})=0\\
\phi\left(x^{2}+y^{2}-u^{2}, (1-u)^{2}-(x+y)^{2}\right)=0."